OpenStudy (anonymous):

a circle is cut in a pack-men shape such that the radius of the sector is x,find the largest possible volume of the cone that is formed from this circle,if the cone has height , h=2x+1

OpenStudy (anonymous):



OpenStudy (phi):

this problem seems to be inconsistent. if x is the radius of the circle then it will correspond to the slant height of the cone and you have x^2 = h^2 + r^2 (r the radius of the cone) but h = 2x+1 and you have x^2 = (2x+1)^2 + r^2 or x^2 = 4x^2 + 4x +1 +r^2 that says the hypotenuse is smaller than one of the sides.... not good.

OpenStudy (anonymous):

i made this munbers by my self ,this problem is my original thought i understand the error but hw can we avoid it how can i change the height

OpenStudy (anonymous):

r=2x+1,h=x

OpenStudy (anonymous):

okay i see the folly thanks guys